OpenStudy (anonymous):

A sphere has a diameter of 18 meters. Find the volume of the sphere, in cubic meters, in terms of π......help me!!!! whats the answerr! >_<

Parth (parthkohli):

The volume of a sphere is \(\Large{4 \over 3}\pi r ^3\)

OpenStudy (anonymous):

4/3Π(r3) is the volume of the sphere. The diameter is 18, so the radius is 9. 4/3Π(93) = 4/3\Pi(729}) = 4/3(2290) = 3053.63